OpenStudy (tyrinzoebrown):

solve each inequality, graph your solution 14. V-6<4 16. N-1/2 ≥ 3/4 18. Q+0.5>-2

OpenStudy (texaschic101):

14. V - 6 < 4 (add 6 to both sides) V < 4 + 6 V < 10 open circle on 10, shading to the left 16. N - 1/2 >= 3/4 (add 1/2 to both sides) N >= 3/4 - 1/2 (common denominator is 4) N >= 3/4 - 2/4 N >= 1/4 closed circle on 1/4, shading to the right 18. Q + 0.5 > -2 (subtract 0.5 from both sides) Q > -2 - 0.5 Q > -2.5 open circle on -2.5, shading to the right

OpenStudy (tyrinzoebrown):

OMGsh thank you so much

OpenStudy (texaschic101):

wait...I messed up on 16..hold on

OpenStudy (tyrinzoebrown):

oh its fine, I got 14 right I just need help with that last 2

OpenStudy (texaschic101):

N - 1/2 >= 3/4 (add 1/2 to both sides) N >= 3/4 + 1/2 (common denominator is 4) N >= 3/4 + 2/4 N >= 5/4 or 1 1/4 closed circle on 5/4 (or 1 1/4), shading to the right

OpenStudy (tyrinzoebrown):

thank you thgank you thank you thank you lol, can you help with 1 more Q though

OpenStudy (texaschic101):

possibly...what you got ?

OpenStudy (tyrinzoebrown):

5x<15 y/3 ≤ 2 25.5q > 102 44.5 <= 2.7d

OpenStudy (tyrinzoebrown):

Qustion is the same

OpenStudy (texaschic101):

5x < 15 (divide both sides by 5) x < 3 open circle on 3, shading to the left y/3 <= 2 (multiply the inequality by 3) y <= 2(3) y <= 6 closed circle on 6, shading to the left 25.5q > 102 (divide both sides by 25.5) q > 102/25.5 q > 4 open circle on 4, shading to the right 44.5 <= 2.7d (divide both sides by 2.7) 44.5/2.7 <= d 16.48 <= d closed circle on 16.48, shading to the right

OpenStudy (tyrinzoebrown):

thanx again

OpenStudy (texaschic101):

no problem :)